OpenStudy (anonymous):

Here are some Math Division Properties of Exponents. Please give me a little help!

OpenStudy (anonymous):

\[\frac{ X^{5/8} }{ X^{3/8} }\]

OpenStudy (anonymous):

\[\frac{ M^{-3} }{ M^{-5 } }\]

OpenStudy (anonymous):

\[\frac{ 3X^{2}Y^{5}Z^{-2} }{ 5XZ^{5} }\]

OpenStudy (anonymous):

\[\left( \frac{ 3X^{2}Y^{5}Z^{-2} }{ 5XZ^{5} } \right)^{3}\]

OpenStudy (anonymous):

\[\left( \frac{ 3X^{2}Y^{5}Z^{-2} }{ 5XZ^{5} } \right)^{-3}\]